# EVA-MILP: Towards Standardized Evaluation of MILP Instance Generation

### v2

@@ 3.3.2 Solving Time Gap
This metric aims to quantify the solver-conditional hardness similarity between the two sets of instances by measuring the percentage difference in their average solving times. A smaller difference indicates higher similarity in hardness for the specific solver used. Prior work, such as [10], has also compared generated and original instances based on relative differences in solving times.
Results The Solving Time Gap (Table 17) reveals varied impacts of generative models on instance hardness. For Set Cover, G2MILP instances showed modest time differences (11–22%) from originals. In Combinatorial Auctions, ACM-MILP dramatically increased solution times (gaps >2400%), while DIG-MILP instances were closer (20–29% gaps). For Independent Set problems, ACM-MILP generated easier instances (times  47–49% shorter). Overall, instance difficulty is highly sensitive to the model and problem type, with mask ratio $\eta$ playing a secondary role. G2MILP (for SC) and DIG-MILP (for CA) better preserved original difficulty levels.
@@ 3.3.3 Hyperparameter Tuning
Automatically optimizing solver hyperparameters is known to be crucial for achieving peak performance on complex algorithms like MILP solvers [31]. Therefore, we employed the Sequential Model-based Algorithm Configuration (SMAC3) framework [32], which utilizes Bayesian optimization. While this optimization approach is general and flexibly extendable to tune various MILP solvers, our current work focused specifically on Gurobi. Following the approach in [13], the tuning process targeted 8 key Gurobi parameters. These parameters were selected as they govern diverse and fundamental components of the MILP solution process, including primal heuristics, search strategy focus, branching decisions, presolving routines, cutting plane generation, and node LP solution methods. The significant impact of these core solver components, and thus the parameters controlling them, on overall performance is well-documented in the MILP literature [e.g., 33, 34, 35]. The objective function for SMAC3 was the minimization of the mean wall-clock solve time across the instances within a designated tuning set. The primary goal of this hyperparameter tuning experiment was to evaluate the generalization capability of the optimized Gurobi configurations using generated MILP instances.
Results  Hyperparameter tuning with SMAC3 (Table 2) yielded varied performance gains. For example, tuning significantly benefited the ACM-MILP (IS, $\eta=0.1$) dataset, reducing solve time by 14.15%. In contrast, improvements for DIG-MILP (CA, $\eta=0.05$) and G2MILP (SC, $\eta=0.1$) were more modest, at 1.15% and 1.59% respectively. These results suggest that the potential for performance enhancement through tuning these specific Gurobi parameters is notably dependent on the dataset and generative model in question.
Table 2: Solving time improvement after Gurobi hyperparameter tuning. Compares the average wall-clock time (seconds) on the test set for the default configuration versus the best configuration found by SMAC3
Dataset Source/Model $\eta$ Default Time (s) Best Time (s) Improvement (%)
IS ACM-MILP 0.05 0.339 0.370 -9.17
@@
0.2 0.339 0.377 -11.13
Original Dataset 0.336 0.366 8.98
CA DIG-MILP 0.01 0.107 0.041 61.35
0.05 0.109 0.042 61.63
0.1 0.108 0.042 61.38
Original Dataset 0.107 0.042 61.26
@@ 3.3.4 Initial Basis Prediction
Following the methodology of [36], we investigated predicting an initial basis for MILP relaxations using a GNN. While adhering to the principles outlined by Fan et al., our implementation was developed independently from scratch. For this purpose, a GNN model was trained specifically on generated MILP instances. The model was trained to predict the basis status (basic or non-basic at lower/upper bounds) for variables and slack variables associated with the MILP relaxation. The GNN’s predictions were subsequently refined into a numerically stable basis using established techniques. To evaluate the effectiveness of training on generated data versus original data for this task, we measured the impact of using the initial basis predicted by the GNN on Gurobi’s performance metrics (solving time, runtime). This was compared against Gurobi’s default setting and potentially a model trained on original data, using unseen test instances. The validity and effectiveness of our re-implemented approach are substantiated by the experimental outcomes.
Results  Table 19 in Appendix.J compares solving performance with GNN-predicted initial basis across different datasets. The effectiveness varied by problem type: For CA problems, both ACM-MILP and DIG-MILP generated instances showed modest improvements (2.6%-4.1%); for IS problems, GNN predictions on generated instances slightly decreased performance (-0.2% to -0.6%); while for SC problems, G2MILP instances showed substantial improvements (9.1%-12.4%), comparable to those on original data. These results indicate that the utility of initial basis prediction is domain-dependent, with Set Cover problems benefiting most significantly.
@@ 3.3.5 Solver-Internal Features
We introduce a novel, computationally efficient methodology to evaluate MILP instance fidelity using solver-internal features: metrics from Gurobi’s deterministic solving process. This approach rapidly assesses and effectively discerns instance set similarities. Its efficacy was validated by a split-half experiment (see Appendix K.1). Figure 2 presents one result of the validation experiments, showing low 1-Wasserstein distances (0.130 in heuristic; 0.188 in root node gap; 0.232 in cut plane) between random original dataset subsets, confirming its capability for authentic inter-set similarity evaluation.
Results Table 3 presents $W_{1}$ distances for root gap distributions (lower $W_{1}$ indicates higher similarity). G2MILP highly replicated SC root gaps (e.g., $W_{1}=0.2294$ at $\eta=0.05$). In contrast, DIG-MILP and ACM-MILP significantly diverged for CA. ACM-MILP showed moderate success for IS at lower $\eta$, diminishing as $\eta$ increased (e.g., $W_{1}\approx 0.87$ for $\eta=0.10$ vs. $W_{1}=1.0388$ for $\eta=0.20$).
Table 3: Comparison of Root Node Gap
Benchmark Generated Dataset Mean Gap (%) Std Dev Gap (%) $W_{1}$ Dist.
Problem Model Ratio ($\eta$) Gen Bench Gen Bench
@@
0.05 7.75 7.61 4.27 4.12 0.2294
0.10 8.05 7.61 4.12 4.12 0.5103
For primal heuristic success distributions (Table 4, $W_{1}$ distance), DIG-MILP (e.g., $W_{1}\approx 0.14$) and ACM-MILP (e.g., $W_{1}\approx 0.34-0.44$) achieved good similarity for CA problems. G2MILP (SC) demonstrated remarkable similarity ($W_{1}\approx 0.08-0.10$). Conversely, ACM-MILP (IS) instances diverged significantly ($W_{1}\approx 17.0-17.2$). Fidelity appears model and problem-dependent, with $\eta$ having a lesser role.
Table 4: Comparison of Heuristic Success Frequency Distributions between Generated (Gen.) and corresponding Benchmark (Base) MILP Instances
Generated Dataset Config Gen. Stats Base Stats Distance
Problem Model $\eta$ Mean Std Dev Mean Std Dev $W_{1}$
@@
IS ACM-MILP 0.05 1.662 0.473 18.883 12.112 17.221
0.10 1.746 0.435 18.883 12.112 17.137
0.20 1.849 0.358 18.883 12.112 17.034
SC G2MILP 0.01 2.619 0.533 2.702 0.491 0.083
0.05 2.600 0.548 2.702 0.491 0.102
0.10 2.598 0.541 2.702 0.491 0.104
For cutting plane usage (Table 5, PCA then $W_{1}$ distance on PC scores), G2MILP (SC) achieved the highest similarity (e.g., PC1 $W_{1}\approx 0.07-0.17$). In contrast, ACM-MILP and DIG-MILP for CA instances showed significant divergences (DIG-MILP PC1 $W_{1}\approx 1.47-1.69$). ACM-MILP (IS) demonstrated intermediate similarity (PC1 $W_{1}\approx 0.89-1.01$). Masking ratio ($\eta$) impact was inconsistent; similarity primarily depends on model-problem combinations.
Table 5: Comparison of Cutting Plane Usage Profiles (1-Wasserstein Distance on Principal Component Scores)
@@
Problem Model Eta ($\eta$) PC1 PC2 PC3
CA ACM-MILP 0.05 1.5570 0.8352 0.5349
0.10 1.4837 0.7155 0.3570
0.20 1.6889 0.3166 0.3731
DIG-MILP 0.01 1.4711 0.6673 0.3601
0.05 1.4868 0.6826 0.3047
0.10 1.5392 0.4394 0.5988
@@
SC G2MILP 0.01 0.0707 0.1271 0.1124
0.05 0.1714 0.1425 0.0677
0.10 0.1374 0.1028 0.1375
Figure 2: Internal Cut-plane Comparison of IS Dataset (Demonstration of Identical Distribution)
@@ 4 Discussion
(i). Superficial structural similarity is an unreliable predictor of computational behavior and difficulty in generated MILP instances. ACM-MILP (CA) instances showed high structural similarity but were exceptionally difficult to solve, with Root Node Gap discrepancies suggesting that structural metrics miss key complexity drivers or that models introduce subtle, difficulty-enhancing variations.
(ii). Simple outcome metrics, such as solving time or branching nodes, do not adequately reveal the source of problem difficulty or underlying changes to instance structure. This is exemplified by ACM-MILP (IS) instances, which, despite some similar outcome metrics (e.g., solving time), exhibited vastly different heuristic behavior, indicating altered internal structures that impact solvability in ways not captured by these surface-level performance measures.
(iii). The effectiveness of generative models is problem-dependent, linked to how well their modification strategies align with the core structural determinants of hardness for specific problem types. G2MILP’s difficulty in generating realistic IS instances is likely due to IS problem hardness being rooted in global graph topology [37, 38], which its local edge constraint modifications struggle to preserve. G2MILP’s better performance on SC problems suggests its direct manipulation of constraint configurations [20] more effectively captures their structural hardness.
While we advocate for the EVA-MILP framework and its emphasis on solver-internal features as a significant advancement, potential limitations warrant acknowledgment. The initial selection of “useful metrics”, though comprehensive, may necessitate ongoing refinement to include emerging or problem-specific indicators and avoid inherent biases. Furthermore, extensive evaluation using rich solver-internal features can be computationally intensive, particularly for large instance sets. Solver-dependency of some internal features also requires careful consideration when generalizing findings. Finally, while EVA-MILP aims to simplify evaluation, comprehensively interpreting multi-faceted data into actionable insights may still demand considerable expertise and remains an area for future development. More limitations are in Appendix. M
@@ 5 Conclusion
The challenges observed in evaluating synthetic MILP instances, particularly the unreliability of traditional structural similarity metrics and the limitations of simple outcome measures, necessitate a paradigm shift in assessment strategies. Our findings consistently demonstrate that generating instances merely possessing superficial structural resemblance does not guarantee they will exhibit comparable computational challenges or elicit solver behaviors akin to real-world problems. We strongly advocate for leveraging solver-internal features as a primary means for a more robust and insightful assessment of instance similarity. In contrast to static features describing only surface ‘appearance’ or outcome metrics offering an aggregate ‘grade’, solver-internal features illuminate the dynamic interplay between an instance’s intrinsic structure and the solver’s algorithmic components. By examining metrics such as cutting plane usage profiles (e.g., the prevalence of Gomory cuts indicating LP relaxation quality [39]) or the success rates of various primal heuristics (suggesting amenability to finding feasible solutions [40]), a more nuanced understanding of the computational challenges inherent in an instance can be achieved.
Focusing on these internal solver dynamics provides a more reliable and fine-grained benchmark. Matching the patterns and distributions of such internal metrics facilitates a comparison based on solver-perceived structure and the actual difficulties encountered during optimization. Our proposed EVA-MILP evaluation framework, which integrates a comprehensive set of useful metrics and supports extensibility, is designed to enable researchers to conveniently assess both generated and non-generated instances. Ultimately, the adoption of such rigorous evaluation methodologies, centered on solver-internal behavior and facilitated by frameworks like EVA-MILP, will be instrumental in developing more representative benchmark instances, thereby fostering more impactful research and development within the field of mathematical optimization.

### v3

@@ 3.3.2 Solving Time Gap
This metric measures the percentage difference in average solving times between generated and original instances, where smaller gaps indicate closer solver-conditional hardness. Table 14 in Appendix shows that G2MILP for SC keeps gaps modest ($\approx$11–22%), DIG-MILP for CA remains near baseline ($\approx$ 20–29%), while ACM-MILP can greatly alter hardness (e.g., 2400%+ increase for CA or $\sim$47–49% faster for IS).
@@ 3.3.3 Hyperparameter Tuning
Automatically optimizing solver hyperparameters is known to be crucial for achieving peak performance on complex algorithms like MILP solvers (Hutter et al., 2011). Therefore, we employed the Sequential Model-based Algorithm Configuration (SMAC3) framework (Lindauer et al., 2022), which utilizes Bayesian optimization. Following the approach in (Liu and others, 2024), the tuning process targeted 8 key Gurobi parameters. These parameters were selected as they govern diverse and fundamental components of the MILP solution process, including primal heuristics, search strategy focus, branching decisions, presolving routines, cutting plane generation, and node LP solution methods. The significant impact of these core solver components, and thus the parameters controlling them, on overall performance is well-documented in the MILP literature (e.g., Lodi and Zarpellon, 2017; Achterberg, 2009; Berthold, 2012). The objective function for SMAC3 was the minimization of the mean wall-clock solve time across the instances within a designated tuning set. The primary goal of this hyperparameter tuning experiment was to evaluate the generalization capability of the optimized Gurobi configurations using generated MILP instances. SMAC3 tuning (Table 3) gives model- and dataset-specific gains. ACM-MILP (IS) decreases about 10% at $\eta=0.1$. DIG-MILP (CA) consistently cuts time by about 61%, and G2MILP (SC) by about 44%, while the original SC set improves only about 3%.
Table 3: Solving time improvement after Gurobi hyperparameter tuning. Compares the average wall-clock time (seconds) on the test set for the default configuration versus the best configuration found by SMAC3.
Dataset Source $\eta$ Default(s) Best(s) Improv. (%)
[…]
@@
[…]
Original 0.336 0.366 8.98
[…]
@@ 3.3.4 Initial Basis Prediction
Following the methodology of (Fan and others, 2023), we investigated predicting an initial basis for MILP relaxations using a GNN. While adhering to the principles outlined by Fan et al., our implementation was developed independently from scratch. For this purpose, a GNN model was trained specifically on generated MILP instances to predict the basis status (basic or non-basic at lower/upper bounds) for variables and slack variables associated with the MILP relaxation. The GNN’s predictions were subsequently refined into a numerically stable basis using established techniques. To evaluate the effectiveness of training on generated data versus original data for this task, we measured the impact of using the initial basis predicted by the GNN on Gurobi’s performance metrics (solving time, runtime). This was compared against Gurobi’s default setting and potentially a model trained on original data, using unseen test instances. The validity and effectiveness of our re-implemented approach are substantiated by the experimental outcomes. Resuls are in Table 4.
Table 4: Impact of Generative Data Augmentation on GNN Initial Basis Prediction. This table evaluates the solving time improvement achieved by initializing Gurobi with a GNN-predicted basis compared to the default initialization. The GNN models were trained either on the Original dataset alone or on the original dataset augmented with synthetic instances generated by different Sources. All performance metrics (average runtime in seconds) are measured on the unseen original test set. Default: Gurobi’s default runtime; Best: Runtime using the GNN-predicted basis; Improv.: Percentage reduction in runtime relative to the default.
Dataset Source $\eta$ Default (s) Best (s) Improv. (%)
CA ACM-MILP 0.05 0.078 0.077 2.6%
0.10 0.079 0.076 2.6%
0.20 0.078 0.075 4.0%
DIG-MILP 0.05 0.079 0.076 3.5%
0.10 0.080 0.076 4.1%
Original 0.078 0.077 2.5%
[…]
@@ 3.3.5 Solver-Internal Features
We introduce a novel, computationally efficient methodology to evaluate MILP instance fidelity using solver-internal features: metrics from Gurobi’s deterministic solving process. We first validated the solver-internal–feature methodology via a split-half experiment (see Appendix K.1). Figure 2 shows one representative result, where random original-set subsets exhibit very low 1-Wasserstein distances (e.g. 0.130 for heuristic, 0.188 for root-node gap, 0.232 for cut-plane profiles), confirming its ability to capture authentic inter-set similarity. To further demonstrate practical viability, we applied the method to the hard ML4CO item_placement dataset under a strict 120-second limit (see more in Section 4.1. This shows that we do not need to fully solve the instance to extract the features. Solver-internal features extracted from Gurobi proved stable and inexpensive to compute. What’s more, comparison with experiments of open-source solvers SCIP and HiGHS (see more in Section 4.2) indicates that, although the framework is general, state-of-the-art solvers like Gurobi provide more precise internal metrics.
Table 5: Comparison of Root Node Gap Distributions
Mean Gap (%) Std Dev (%)
Problem Model $\eta$ Gen Original Gen Original $W_{1}$ Dist.
@@
[…]
Table 6: Comparison of Heuristic Success Frequency Distributions
Gen. Stats Base Stats
Problem Model $\eta$ Mean Std Mean Std $W_{1}$ Dist.
@@
[…]
Results Table 5 presents $W_{1}$ distances for root gap distributions (lower $W_{1}$ indicates higher similarity). G2MILP highly replicated SC root gaps (e.g., $W_{1}=0.2294$ at $\eta=0.05$). In contrast, DIG-MILP and ACM-MILP significantly diverged for CA. ACM-MILP showed moderate success for IS at lower $\eta$, diminishing as $\eta$ increased.
Table 7: Comparison of Cutting Plane Usage Distributions
@@
Problem Model $\eta$ PC1 PC2 PC3
[…]
$0 \cellcolorgray!2$01.6889 0.3166 0.3731
[…]
@@
[…]
For primal heuristic success distributions (Table 6, $W_{1}$ distance), DIG-MILP (e.g., $W_{1}\approx 0.14$) and ACM-MILP (e.g., $W_{1}\approx 0.34-0.44$) achieved good similarity for CA problems. G2MILP (SC) demonstrated remarkable similarity ($W_{1}\approx 0.08-0.10$). Conversely, ACM-MILP (IS) instances diverged significantly ($W_{1}\approx 17.0-17.2$).
For cutting plane usage (Table 7), G2MILP (SC) achieved the highest similarity (e.g., PC1 $W_{1}\approx 0.07-0.17$). In contrast, ACM-MILP and DIG-MILP for CA instances showed significant divergences (DIG-MILP PC1 $W_{1}\approx 1.47-1.69$). ACM-MILP (IS) demonstrated intermediate similarity (PC1 $W_{1}\approx 0.89-1.01$).
Figure 2: Internal Cut-plane Comparison of IS (Demonstration of Identical Distribution)
4 Experiments
4.1 Efficiency Test on Super Hard Instances
To validate the practical feasibility and robustness of our solver-internal feature analysis method, we designed an efficiency test. The core goal of this experiment is to demonstrate that even under strict computational constraints (i.e., tight time limits) and when facing a well-known set of hard instances, our method can still stably and efficiently extract meaningful feature distributions. This supports our key hypothesis that obtaining such deep behavioral features does not require solving instances to optimality, thereby making the evaluation both low-cost and highly efficient.
Table 8: W-1 distances of solver-internal features across item placement halves
Feature W-1 Dist.
Root Node Gap 0.7613
Heuristic Success Count 0.7320
Cut Plane Usage
   – PC1 0.0917
   – PC2 0.0870
   – PC3 0.0904
Dataset Selection: We used the item placement dataset from the ML4CO competition, which is recognized for its high solving difficulty. A total of 1000 instances were selected for testing.
Computational Budget Constraint: To simulate scenarios with limited computational resources, we imposed a strict time limit of 120 seconds for the Gurobi solver on each instance.
Validation Method: We employed a split-half validation approach to evaluate the stability of the extracted features. The 1000 solver logs were randomly divided into two halves of 500 instances each. We then independently extracted solver-internal features—including Root Node Gap, Heuristic Success Count, and Cut Plane Usage—from both halves and computed the distributional similarity between them. Strong stability and reliability of the extraction method are indicated if these random subsets, originating from the same underlying distribution, exhibit highly similar feature distributions (i.e., a low Wasserstein distance).
The split-half validation demonstrates remarkable stability in the extracted solver-internal features, even under the strict 120-second time limit. As shown in Table 8, the Cut Plane Usage metric exhibits exceptionally low 1-Wasserstein distances for its top three principal components (PC1: 0.0917, PC2: 0.0870, and PC3: 0.0904). These near-zero values indicate that the solver’s strategic application of cuts—a crucial early-stage behavior—is highly consistent across the two random data subsets. Similarly, the Root Node Gap (0.7613) and Heuristic Success Count (0.7320) show strong distributional similarity, confirming that a stable signal for these metrics can also be captured long before the solver reaches optimality. Collectively, these results validate that a truncated solving process is sufficient to extract a reliable and consistent ”fingerprint” of solver behavior, confirming the low-cost and efficient nature of our methodology.
Conclusion: This experiment strongly demonstrates the practical value of our framework: 1 High Efficiency and Low Cost: The results confirm that a truncated solving process (120-second timeout) is sufficient to extract stable and representative solver-internal feature distributions. This substantially reduces the time and computational cost of evaluating large-scale or extremely hard instance sets. 2 Feature Stability: The very small Wasserstein distances observed for Root Node Gap and Cut Plane Usage indicate that these early-stage solver behavior features are highly stable. 3 Practical Feasibility: The success of this test shows that the GenBench-MILP framework is not only theoretically sound but also practically applicable for evaluating MILP instances that are difficult to solve to optimality within a reasonable time frame and that reflect real-world complexity.
4.2 Cross-Solver Comparison and Stability Analysis
To assess the generality of the GenBench-MILP framework and probe the internal stability of different solvers, we extended our evaluation to include the open-source solvers SCIP and HiGHS. We conducted the same IS split-half validation procedure, using Gurobi as the stable baseline for comparison. Table 9 summarizes the key results.
Table 9: 1-Wasserstein distances across solvers on IS split-half validation
Feature Gurobi (Baseline) SCIP HiGHS
[…]
The data reveals a complex and multi-faceted picture of solver stability. First, examining the Root Node Gap, we observe dramatic differences. Gurobi exhibits exceptional stability with a Wasserstein distance of just 0.1884. In stark contrast, HiGHS shows a distance of 26.7297, over 140 times larger than Gurobi’s, signifying severe instability in its root node analysis. This suggests that the initial problem assessment and relaxation strategies in HiGHS are profoundly sensitive to small changes in problem structure. Second, the results for the primary principal component of Heuristic features show a similar trend, though less pronounced. Gurobi remains the most stable (0.1300), while SCIP and HiGHS are approximately 2.5x and 1.8x less stable, respectively. This implies Gurobi’s primal heuristic strategies are more robust. Interestingly, the analysis of Cut Plane features presents a contrasting result. Here, HiGHS is by far the most stable solver (0.0612), with a consistency nearly 4 times greater than Gurobi’s (0.2318). This suggests that HiGHS may employ a more deterministic or conservative cut generation strategy, whereas Gurobi and SCIP might use more adaptive, opportunistic approaches that are inherently more variable.
From this detailed analysis, we draw two main conclusions. First, the successful application to SCIP and HiGHS validates that the GenBench-MILP framework is indeed solver-agnostic and generalizable. Second, and more importantly, the quality of instance evaluation is intrinsically linked to solver stability, which is not monolithic but highly feature-dependent. Gurobi’s superior stability in the critical root-node phase makes it the most reliable choice for a baseline. The extreme volatility of the root node gap in HiGHS, for instance, could obscure subtle computational similarities that a more stable solver like Gurobi can reliably detect.
In summary, this cross-solver comparison highlights a powerful secondary application of our framework. Beyond its primary purpose of instance evaluation, GenBench-MILP serves as an effective diagnostic tool for profiling the internal dynamics of MILP solvers, offering quantitative insights into their stability and behavioral responses to perturbations across different strategic components.
@@ 5 Discussion
Finding 1: Superficial structural similarity is an unreliable predictor of computational behavior and difficulty in generated MILP instances. ACM-MILP (CA) instances showed high structural similarity but were exceptionally difficult to solve, with Root Node Gap discrepancies suggesting that structural metrics miss key complexity drivers or that models introduce subtle, difficulty-enhancing variations.
Finding 2: Simple outcome metrics, such as solving time or branching nodes, do not adequately reveal the source of problem difficulty or underlying changes to instance structure. This is exemplified by ACM-MILP (IS) instances, which, despite some similar outcome metrics (e.g., solving time), exhibited vastly different heuristic behavior, indicating altered internal structures that impact solvability in ways not captured by these surface-level performance measures.
Finding 3: The effectiveness of generative models is problem-dependent, linked to how well their modification strategies align with the core structural determinants of hardness for specific problem types. G2MILP’s difficulty in generating realistic IS instances is likely due to IS problem hardness being rooted in global graph topology (Tarjan and Trojanowski, 1977; West, 2001), which its local edge constraint modifications struggle to preserve. G2MILP’s better performance on SC problems suggests its direct manipulation of constraint configurations (Balas and Ho, 1980) more effectively captures their structural hardness.
@@ 6 Conclusion
We introduce GenBench-MILP , a comprehensive benchmark framework designed to address the critical shortcomings of traditional evaluation methods for generated MILP instances. Our findings reveal that metrics based on superficial structural similarity or simple outcomes often fail to capture an instance’s true computational nature and difficulty. GenBench-MILP facilitates a more accurate and holistic assessment by shifting the focus towards solver-perceived difficulty and the practical utility of instances in downstream applications. By enabling this multifaceted methodology, our framework will be instrumental in guiding the development of higher-quality instance generators and fostering impactful research within the field of mathematical optimization.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.